William is 5 years older than Roberto. Roberto is 3 years older than Ashley. If Ashley is 14 years old, how old is William?
Answer by 932982(3)

William= 5+ Robert
Robert= 3+ Ashley
Ashley= 14
Robert= 3+14(Ashley's age) = 17
William= 5+17(Robert's age) =22
William= 22 years old